4x-6y=5
6x-9y=4
x= 3.5
y= 1.5
ans:3.5,1.5

Found 2 solutions by stanbon, Adittadas4@gmail.com:
Answer by stanbon(75887)

4x-6y=5
6x-9y=4
----
Multiply thru the top equation by -3
Multiply thru the bottom equation by 2
-12x + 18y = -15
12x - 18y = 8
-----
Adding you get:
0 + 0 = -7
That is a contradiction and indicates
there is no solution for the systerm.
----
Note: If you graped the two equations you
would find that they are parallel.
Cheers,
Stan H.

Answer by Adittadas4@gmail.com(2)

Ans:x=3.5
y=1.5